• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name pking

    Log entry time 18:58:00 on November14,2009

    Entry number 299880

    keyword=Target: Looks fine at normal beam current.

    Request 25 uA on the LD2 target, then 100 uA.

    All target parameters seem to behave normally. Back to production!

    FIGURE 1

    FIGURE 2